OpenStudy (anonymous):

Gold in the form of solid cylinders will be molded into solid blocks. Each cylinder has a diameter of 6.19 cm and a height of 6 cm. Each solid block will have a volume of 360 cm3. How many cylinders of gold will it take to mold one block of gold?

OpenStudy (lucaz):

assuming pi = 3.14 vol. cylinder = area of base * height v = pi*3.095^2 * 6 = approx. 180.47 cm^3 2 cylinders = 2*180 = 360 cm^3 it will take two cylinders to mold one block

OpenStudy (anonymous):

Thanks :)

OpenStudy (lucaz):

no problem